# SMiR: Efficient Synthetic Data Pipeline to Improve Multi-image Reasoning

### v1

@@ Abstract
Vision-Language Models (VLMs) have shown strong performance in understanding single images, aided by numerous high-quality instruction datasets. However, multi-image reasoning tasks are still under-explored in the open-source community due to two main challenges: (1) scaling datasets with multiple correlated images and complex reasoning instructions is resource-intensive and maintaining quality is difficult, and (2) there is a lack of robust evaluation benchmarks for multi-image tasks. To address these issues, we introduce SMiR, an efficient synthetic data-generation pipeline for multi-image reasoning, and a high-quality dataset generated using this pipeline. Our pipeline efficiently extracts highly correlated images using multimodal embeddings, combining visual and descriptive information and leverages open-source LLMs to generate quality instructions. Using this pipeline, we generated 160K synthetic training samples, offering a cost-effective alternative to expensive closed-source solutions. Additionally, we present SMiR-Bench, a novel multi-image reasoning evaluation benchmark comprising 200 diverse examples across 7 complex multi-image reasoning tasks. SMiR-Bench is multi-turn and utilizes a VLM judge to evaluate free-form responses, providing a comprehensive assessment of model expressiveness and reasoning capability across modalities. We demonstrate the effectiveness of SMiR dataset by fine-tuning several open-source VLMs and evaluating their performance on SMiR-Bench. Our results show that models trained on our dataset outperform baseline models in multi-image reasoning tasks up to 8% with a much more scalable data pipeline. Our codebase is available at https://github.com/togethercomputer/SMiR.
@@ 1 Introduction
Vision-Language Models (VLMs) have shown impressive capabilities in tasks involving single images, particularly open-source models that have benefited from high-quality instruction datasets (Laurençon et al., 2024b; Zhang et al., 2023; Xu et al., 2022). However, in multi-image tasks, such as comparing or analyzing relationships between multiple images, the performance of open-source VLMs (Liu et al., 2024b; a; Li et al., 2024a; Awadalla et al., 2023; Yao et al., 2024; Wang et al., 2023b) significantly lags behind that of their closed-source counterparts like GPT-4 (Achiam et al., 2023), Claude 3.5 Sonnet (Anthropic, 2024a), Claude 3 (Anthropic, 2024b), and Gemini 1.5 (Reid et al., 2024).
While we do not have full visibility into how open-source models are trained, a reasonable assumption for their underperformance in multi-image reasoning tasks is the advantage closed-source models have in terms of training datasets. Closed-source developers likely benefit from access to large, high-quality datasets, which are critical for training advanced models capable of reasoning across multiple images. In contrast, open-source models may face limitations in accessing or constructing datasets that are specifically tailored to these tasks, which impacts their overall performance.
One major challenge lies in collecting and curating large-scale datasets with highly correlated images. Identifying meaningful connections between images requires vast amounts of data and sophisticated algorithms, which most multi-image instruction-tuning datasets lack. For example, MANTIS (Jiang et al., 2024) often pairs unrelated images in the same reasoning task, unintentionally simplifying the challenge. Consider a question like: Compare the behaviors and characteristics shown in these images. How do they demonstrate different aspects of survival and adaptation? Previous methods may match it with a relevant image of a dog alongside completely unrelated images like a drone hovering in the air and shipping containers at a port. While the question naturally applies to analyzing an animal’s behavioral adaptations, its pairing with technological and industrial objects (drones, logistics, etc.) forces superficial comparisons that dilute the biological focus of the inquiry. This makes the task easier, as the unrelated items can only be discussed through loose metaphorical connections to survival and adaptation.
Figure 1: Comparison of top (example of MANTIS dataset sharegpt4v-pt_125, which can have a data sample with animals, but are concatenated together with unrelated images) vs. bottom (SMiR dataset where related images are together) for multi-image reasoning.
In contrast, comparing related images—like a series of different animals in their natural habitats—demands a deeper understanding of predator behavior and ecological adaptation. Here, the task requires a more nuanced comparison of hunting techniques, territorial behaviors, and social structures within the same family of species. MMDU-45K (Liu et al., 2024e) attempts to address this by clustering image captions and category tags, but it overlooks the crucial visual relationships. These shortcomings highlight the need for datasets with truly correlated images to foster complex multi-image reasoning tasks that better reflect real-world analytical challenges.
Second, scaling up the number of highly correlated images presents significant challenges. Existing datasets such as MANTIS, MMDU-45K, and MultiInstruct (Xu et al., 2022) require extensive human curation and annotation, resulting in a labor-intensive and time-consuming process. To minimize human effort, researchers have leveraged GPT-4 family models (Peng et al., 2023; Wang et al., 2023a) to generate synthetic datasets—including MMInstruct (Liu et al., 2024d), Multimodal ArXiv (Li et al., 2024b), MIMIC-IT (Li et al., 2023a), StableLLaVA-Instruct (Li et al., 2023c), and SVIT-Instruct (Zhao et al., 2023). However, this method proves expensive and difficult to scale effectively.
Third, evaluating multi-image reasoning is challenging due to the limitations of traditional assessment methods. Multiple-choice formats, used in previous multi-image reasoning benchmarks (Meng et al., 2024; Fu et al., 2024; Wang et al., 2024a; Yue et al., 2024), fail to capture the nuanced reasoning process and justifications behind a model’s answers. This is problematic because models can achieve similar performance on multiple-choice tasks while having vastly different reasoning capabilities. In contrast, free-form responses allow models to demonstrate their thought processes, providing a more comprehensive evaluation of their true reasoning abilities and distinguishing between models that rely on superficial pattern matching and those with a deeper understanding.
To address these challenges, we propose a synthetic data generation pipeline, SMiR, for multi-image reasoning, along with a human-annotated evaluation benchmark, SMiR-Bench. SMiR generates correlated and challenging multi-image reasoning questions, while SMiR-Bench evaluates models using free-response tasks in complex multi-image scenarios.
Our contributions are as follows:
  • We present SMiR, a pipeline for multi-image instruction tuning. To construct SMiR, we apply existing algorithms in a novel way to collect multiple images with high correlation. Our approach combines cluster sampling for data quality robustness and graph iteration sampling for diversity. Both algorithms leverage multimodal embeddings that integrate image and caption information to effectively group correlated images, enabling challenging multi-image instruction tuning scenarios.
@@
  • We develop a scalable synthetic multimodal data generation pipeline that leverages open-source LLMs such as Meta Llama 3.1 70B Instruct Turbo (Dubey et al., 2024). This approach eliminates the need for expensive closed-source models, reducing costs by up to 50 times and increasing speed by up to 10 times (Kirkovska, 2024), while significantly minimizing human annotation efforts.
  • We present a new multimodal evaluation benchmark that uses multi-turn interactions and free-form responses to assess both final answers and reasoning processes in complex multi-image tasks. The evaluation is conducted pairwise, directly comparing model performances. Using reference answers from GPT-4-Turbo and other open-source models as baselines, we observe up to an 8% improvement in win rate for instruction-tuned models trained on the SMiR dataset.
2 Related Works
Vision Language Models
We focus on instruction tuning Vision-Language Models (VLMs) that utilize a pretrained Large Language Model (LLM) backbone because this approach is cost-effective and more accessible for the open-source community. Since the backbone responsible for language understanding is already trained, the overall training process becomes simpler and requires fewer resources. Our primary task involves aligning the vision encoder—typically architectures like Vision Transformer (ViT) (Dosovitskiy, 2020), SigLIP (Zhai et al., 2023), or CLIP (Radford et al., 2021)—with the LLM backbone. This alignment is facilitated through linear layers that connect the vision encoder to the backbone, enabling the integration of visual and textual information. For instance, BLIP-2 (Li et al., 2023b) uses OPT (Zhang et al., 2022) and FLAN-T5 (Chung et al., 2022) as backbones, MiniGPT-4 (Zhu et al., 2023) utilizes Vicuna (Chiang et al., 2023), and Qwen-2-VL (Wang et al., 2024b) employs Qwen-2-1.5B (Yang et al., 2024) as the language backbone. In this paper, we focus on creating a high-quality multi-image reasoning dataset for instruction tuning instead of large-scale interleaved pretraining datasets like OBELICS (Laurençon et al., 2024a), MINT-1T (Awadalla et al., 2024), and LAION-5B (Schuhmann et al., 2022).
Multi-Image Reasoning Data
Recent advancements in multi-image reasoning instruction tuning datasets include MANTIS (Jiang et al., 2024) and MMDU-45K (Liu et al., 2024e), both aiming to improve reasoning capabilities in VLMs. However, these datasets have limitations in their approaches. MANTIS randomly concatenates single image pairs from LLaVA-665k (Liu et al., 2024a), which often results in uncorrelated images within multi-image scenarios, potentially undermining the complexity of reasoning tasks. MMDU-45K attempts to address this issue by utilizing sentence transformers (Reimers, 2019) with description text and clustering techniques to group related images, but does not consider visual components. The dataset is then further enhanced, assisted by GPT-4 to generate comprehensive answers for the grouped images. Building upon these efforts, SMiR introduces a novel approach that leverages both vital visual and caption information to ensure highly correlated images within multi-image sets with the use of open-source LLMs. These scalable methods leads to the generation of more challenging questions that require deeper analysis and understanding of visual relationships, pushing the boundaries of multi-image reasoning capabilities in VLMs.
@@ 2 Related Works
Multi-Image Reasoning Benchmarks
Recent VLM benchmarks (Chiang et al., 2024; Lin et al., 2024; Liu et al., 2024c) have made strides by incorporating free-response evaluations, marking a significant improvement over traditional multiple-choice formats. However, these benchmarks still lack a comprehensive approach that combines automatic, multi-turn, and pairwise evaluation capabilities. Our benchmark addresses this gap, drawing inspiration from Auto-Hard-Auto v0.1 (Li et al., 2024c). We have adapted and expanded this framework to enable robust multimodal evaluation, providing a more holistic assessment of VLM performance across complex, multi-image reasoning tasks. This approach allows for a deeper analysis of both the final answers and the underlying reasoning processes employed by VLMs in real-world SMiR-Bench scenarios.
Figure 2: Our end-to-end pipeline: from sampling and LLM prompting to conversation generation. The example conversation is based on a sports scenario, demonstrating the pipeline’s ability to generate contextually relevant multi-turn dialogues.
@@ 3 SMiR: Synthetic Multi-Image Reasoning Data Pipeline
SMiR pipeline efficiently generates complex multi-image reasoning data through a carefully designed methodology that reflects key design considerations and trade-offs. Starting with a large-scale image-caption dataset $D$ containing $N$ image-caption pairs $(I_{i},C_{i})_{i=1}^{N}$, SMiR first constructs multimodal embeddings for each pair. Then, grouping algorithms identify correlations between these embeddings. Finally, open-source LLMs are used to generate complex question-answer pairs based on the sampled correlations. In the following subsections, we describe our approach’s core components—the construction of multimodal embeddings, grouping algorithms, and synthetic data generation implementation—while examining the critical design decisions behind our sampling algorithms, prompts, and data sources that shaped the final SMiR dataset.
@@ 3.1 Multimodal Embedding Construction
While embedding-based approaches have been used before to capture relationships between images and text, relying solely on either image or caption embeddings limits our ability to fully understand the connections between visual content and descriptive text. A more comprehensive approach requires incorporating both contexts simultaneously.
To address this, we developed a method that combines image and caption information into a single multimodal embedding. Our approach utilizes SigLIP or CLIP embeddings for the images, combined with corresponding caption embeddings. We formulated the multimodal embedding as follows:
$E_{multimodal}=E_{image}+c\cdot E_{caption}$ (1)
where $E_{multimodal}$ is the multimodal embedding, $E_{image}$ is the image embedding, and $E_{caption}$ is the caption embedding. The constant $c$ adjusts the weight of the caption embedding. This value is determined empirically through human evaluation of a small sample of generated questions. For the ShareGPT4V dataset (Chen et al., 2023), $c=0.2$ worked well, though this parameter may vary for different datasets depending on their quality and the descriptiveness of their captions to achieve optimal results.
This multimodal approach allows us to capture nuanced relationships between visual content and descriptive text, enhancing our ability to identify and group correlated images effectively. After constructing the multimodal embeddings, we applied UMAP (McInnes et al., 2018) to reduce the dimensionality of the vectors, projecting them into a lower-dimensional space for more efficient analysis and visualization, while preserving the essential structure of the data.
3.2 Grouping Images
Figure 3: Left: Cluster Matching from Different Embedding Spaces, where images are sampled only from the union of clusters with the same color. Right: Sampling based on embedding distance for increased diversity from initial image (orange), where all images are considered within a single question.
Greedy Cluster Matching
1) Orders clusters by size in descending order
This process continues until all clusters are matched or one set is exhausted, demonstrating high efficacy in producing quality image-caption pairs through dual-embedding validation. The score function, which measures overlap normalized by average cluster size, ensures each cluster corresponds to a semantically similar cluster from the other model. The detailed steps of this algorithm and pseudocode are presented in Figure 3 and Algorithm 1 (Appendix A.1).
While this approach effectively leverages both embedding models to confirm spatial relationships and associated semantic meanings, with images matched within clusters being corroborated by two independent embedding models, it has a notable limitation: sampling confined to matched clusters can lead to overly specialized image subjects, such as clustering only sheep-related images rather than diverse animal scenes, as demonstrated in Figure 6 (Appendix B.1). This specialization occurs because sampling is restricted to a single matched cluster, potentially limiting the diversity of selected images.
Random Sampling with Iteration
To address the limitations of cluster-confined sampling and transcend individual cluster boundaries, we developed Random Sampling with Iteration as our preferred solution. This method generates more generalized image subjects while maintaining semantic relationships through an iterative sampling approach. Given a set of embeddings $X=\{x_{1},...,x_{n}\}$, the algorithm:
1) Randomly selects an initial embedding index and initializes the selected set $S$
2) For each subsequent selection until reaching $N$ samples, computes a probability distribution over remaining embeddings based on: $p(x_{j})\propto\frac{1}{\sum_{u\in S}\|x_{j}-x_{u}\|^{k}+\epsilon}$, where $x_{j}\in X\setminus S$
@@ 3.2 Grouping Images
3) Samples the next embedding according to this distribution, favoring points with smaller cumulative distances
Here, $k$ (default: 12) controls the emphasis on proximity - higher values more strongly favor nearby points by amplifying the differences in the distance calculations. The flexibility of this random vector sampling approach enables drawing from semantically related regions of the embedding space, while maintaining some randomness through probabilistic selection. This contributes to a more coherent dataset while still allowing for sufficient variety in the samples. For example, as shown in Figure 7 (Appendix B.2), this approach successfully samples related animal images while maintaining semantic coherence, in contrast to the disconnected clusters seen earlier. The detailed steps of this algorithm and pseudocode are presented in Figure 3 and Algorithm 2 (Appendix A.2).
@@
By assembling related images before prompting the LLM, we create a more coherent and contextually rich input, enabling the model to generate more nuanced and relevant synthetic data for multi-image reasoning tasks.
3.3 Implementation
Our implementation of the SMiR dataset construction builds upon this approach of leveraging grouped images. After sampling correlated image-caption pairs, we developed a synthetic data generation process that incorporates these captions into system prompts for an open-source LLM. We specifically chose Meta Llama 3.1 70B Turbo for its significant efficiency advantages—up to 50 times cheaper and 10 times faster compared to GPT-4 (Kirkovska, 2024). This process generates complex multi-turn conversations between User and Assistant as seen in Table 2, with questions tailored to the selected images, as shown in Figure 6 and Figure 7 (Appendix B).
Our implementation employs two distinct types of system prompts to effectively capture different aspects of visual reasoning. The first type focuses on shorter visual questions, often involving OCR tasks, while the second addresses longer reasoning questions requiring in-depth analysis. To effectively capture the nuances of both complex reasoning scenarios and straightforward visual comprehension tasks, we developed separate, tailored prompts for each question type. These prompts were inspired by the methodologies of LLaVA (Liu et al., 2024b) and Chain-of-Thought (Wei et al., 2022), designed to generate multi-turn conversations that enhance the complexity and depth of interactions in our dataset. The detailed implementations of short-form prompts for visual tasks and long-form prompts for extended reasoning are presented in Appendices C.1 and C.2, respectively.
For the data source, we leveraged ShareGPT4V (Chen et al., 2023), an open-source dataset that combines data from LLaVA-Instruct (Liu et al., 2024a) and COCO (Lin et al., 2014). ShareGPT4V was initially developed using 100,000 captions generated by GPT4-Vision and subsequently expanded to 1.2 million image-caption pairs through an improved caption model, maintaining both quality and scale throughout the expansion process. To ensure diversity in our synthetic data, we processed the dataset in batches of 20,000 images, generating 5,000 synthetic conversations per batch. Each synthetic conversation contains an average of 4.65 images, and in total, we used 640,000 image-caption pairs to produce 160,000 synthetic conversations. We employed a vector sampling algorithm in conjunction with our long question prompts, with the primary objective of generating questions that were not only diverse but also challenged multi-step reasoning capabilities through multi-turn interactions. The modular design of our implementation allows for easy adaptation to other data resources in future applications.
Table 2: SMiR Dataset Statistics

### v2

@@ Abstract
Vision-Language Models (VLMs) excel at understanding single images, aided by high-quality instruction datasets. However, multi-image reasoning remains underexplored in the open-source community due to two key challenges: (1) scaling datasets with correlated images and complex reasoning instructions is resource-intensive, and (2) robust evaluation benchmarks for multi-image tasks are lacking. To address this, we introduce SMiR, a synthetic data-generation pipeline for multi-image reasoning, along with a high-quality dataset generated using this pipeline. SMiR efficiently extracts correlated images via multimodal embeddings, integrates visual and descriptive information, and leverages open-source LLMs to generate quality instructions. Using this approach, we produce 160K synthetic training samples, offering a cost-effective alternative to closed-source solutions. Additionally, we present SMiR-Bench, a multi-image reasoning benchmark comprising 200 diverse examples across seven complex reasoning tasks. SMiR-Bench is multi-turn and employs a VLM judge to evaluate free-form responses, providing a comprehensive assessment of model expressiveness and reasoning capability across modalities. We demonstrate the effectiveness of SMiR by fine-tuning open-source VLMs and evaluating them on SMiR-Bench. Our codebase is available at https://github.com/togethercomputer/SMiR.
@@ 1 Introduction
Vision-Language Models (VLMs) have demonstrated strong performance on single-image tasks, particularly open-source models trained with high-quality instruction datasets (Laurençon et al., 2024b; Zhang et al., 2023; Xu et al., 2022). However, open-source VLMs struggle with multi-image reasoning tasks—such as comparing and analyzing relationships between multiple images—lagging behind closed-source models like GPT-4 (Achiam et al., 2023), Claude 3.5 Sonnet (Anthropic, 2024), and Gemini 1.5 (Reid et al., 2024). This performance gap likely stems from the superior multi-image datasets available to closed-source models, which are essential for effective reasoning across multiple images.
A key challenge in multi-image reasoning is collecting and curating large-scale datasets with highly correlated images. Existing datasets such as MANTIS (Jiang et al., 2024) often pair unrelated images, simplifying the task and limiting the complexity of reasoning. MMDU-45K (Liu et al., 2024f) improves upon this by clustering images based on captions and category tags but overlooks purely visual relationships. Consequently, datasets that explicitly ensure image correlation are needed to advance complex multi-image reasoning.
Scaling such datasets further exacerbates the challenge. Human curation, as required in datasets like MultiInstruct (Xu et al., 2022), is resource-intensive. While synthetic datasets generated via GPT-4 (Peng et al., 2023; Wang et al., 2023) offer a scalable alternative (Liu et al., 2024e; Li et al., 2024a; 2023a), they remain expensive and difficult to scale effectively.
Another challenge is evaluation. Existing multi-image benchmarks, such as MMIU (Meng et al., 2024) and MMMU (Yue et al., 2024), rely on multiple-choice questions, failing to assess the depth of a model’s reasoning. While some benchmarks incorporate free-form responses (Liu et al., 2024c), they lack multi-turn interactions, which are essential for evaluating complex reasoning processes.
To address these issues, we introduce SMiR, a synthetic data generation pipeline for multi-image reasoning, and SMiR-Bench, a multi-image evaluation benchmark. Our key contributions are:
  • We propose SMiR, a pipeline that efficiently identifies correlated images using multimodal embeddings, applying cluster sampling and graph iteration sampling to ensure diversity and robustness in instruction tuning.
@@
  • We develop a scalable synthetic data generation pipeline using open-source LLMs such as Meta Llama 3.1 70B Instruct Turbo (Dubey et al., 2024), reducing costs by up to 50x and increasing speed by up to 10x compared to closed-source alternatives (Kirkovska, 2024).
  • We introduce SMiR-Bench, a multi-turn evaluation benchmark that assesses models via free-form responses and reasoning justifications, using GPT-4-Turbo and other baselines for comparison. Models trained on SMiR improve win rates by up to 8% over existing instruction-tuned models.
By advancing both dataset generation and evaluation, SMiR facilitates the development of stronger open-source VLMs for multi-image reasoning.
@@
Figure 1: Our end-to-end pipeline converts image-caption pairs into synthetic multi-turn conversations using multimodal embeddings, strategic sampling, and LLM prompting. The example, based on a sports scenario, illustrates how the pipeline generates contextually rich dialogues by leveraging visual relationships.
Figure 2: Comparison of an example from the MANTIS dataset, which includes samples from ShareGPT4V-PT where unrelated images such as animals and other objects are concatenated (top), vs. an example from the SMiR dataset, where only related images are grouped together (bottom), for multi-image reasoning.
@@ 2 SMiR: Synthetic Multi-Image Reasoning Data Pipeline
Starting with a large-scale image-caption dataset $D$ containing $N$ image-caption pairs $(I_{i},C_{i})_{i=1}^{N}$, SMiR first constructs multimodal embeddings for each pair. Then, sampling algorithms identify correlations between these embeddings. Finally, open-source LLMs are used to generate complex question-answer pairs based on the sampled correlations. The end-to-end pipeline is shown in Fig. 1.
@@ 2.1 Multimodal Embedding Construction
Embedding-based approaches have been widely used to model relationships between images and text. However, relying solely on image or caption embeddings limits the ability to fully capture the connections between visual content and descriptive text. A more comprehensive approach requires integrating both modalities.
To achieve this, we construct a multimodal embedding by combining SigLIP or CLIP image embeddings with their corresponding caption embeddings:
$E_{\text{multimodal}}=E_{\text{image}}+c\cdot E_{\text{caption}}$ (1)
where $E_{\text{multimodal}}$ represents the fused embedding, and $c$ adjusts the weight of the caption embedding. This parameter is tuned empirically based on small-scale human evaluations. For the ShareGPT4V dataset (Chen et al., 2023), $c=0.2$ provided the best balance between visual and textual contributions. The optimal value may vary depending on dataset quality and caption descriptiveness.
This multimodal approach enables better identification and clustering of correlated images by capturing both visual and textual relationships. We apply UMAP (McInnes et al., 2018) for dimensionality reduction, preserving essential data structures while facilitating efficient analysis and visualization.
2.2 Random Sampling with Iteration
After constructing multimodal embeddings, we apply a random sampling algorithm to select semantically related images, ensuring diversity while preserving coherence. Given a set of embeddings $X=\{x_{1},...,x_{n}\}$, the algorithm proceeds as follows:
  1. 1.
    Randomly selects an initial embedding and initializes the selected set $S$.
  2. 2.
    Iteratively selects embeddings until reaching $N$ samples, using the probability distribution:
    $p(x_{j})\propto\frac{1}{\sum_{u\in S}\|x_{j}-x_{u}\|^{k}+\epsilon},\quad x_{j}\in X\setminus S$ (2)
  3. 3.
@@ 2.2 Random Sampling with Iteration
    Samples the next embedding based on this probability, favoring closer points.
Here, $k$ (default: 12) controls the emphasis on proximity—higher values prioritize nearby points more strongly by amplifying distance differences. This approach balances semantic coherence and diversity, allowing for meaningful multi-image relationships while introducing some randomness.
For example, as shown in Fig. 6 (Section C.2), this method effectively clusters related animal images while avoiding the disconnected groupings seen in simpler sampling methods. The full algorithm and pseudocode are provided in Algorithm 2 (Section B.2).
@@
By assembling related images before prompting the LLM, we generate richer contextual inputs, enabling the model to produce more nuanced and relevant synthetic data for multi-image reasoning tasks.
Table 2: Statistics of the SMiR dataset
@@
2.3 Implementation
Our implementation of the SMiR dataset construction builds on the structured sampling of correlated image-caption pairs. After grouping related images, we generate synthetic data by incorporating these captions into system prompts for an open-source LLM. We use Meta Llama 3.1 70B Turbo due to its efficiency, as it is up to 50 times cheaper and 10 times faster than GPT-4 (Kirkovska, 2024). This process produces complex, multi-turn conversations between the user and assistant, as shown in Table 2. Sampled images and corresponding generated questions are illustrated in Fig. 5 and Fig. 6 (Appendix C).
To capture different aspects of visual reasoning, we design two types of system prompts. The first type focuses on shorter visual questions, often requiring OCR-based understanding. The second type is designed for long-form reasoning, where multi-step inference and deeper contextual understanding are needed. These prompts are inspired by LLaVA (Liu et al., 2024b) and Chain-of-Thought reasoning (Wei et al., 2022), enabling the generation of multi-turn conversations that enhance dataset complexity. The detailed structures of these prompts are provided in Section D.1 and Section D.2.
For data sources, we use ShareGPT4V (Chen et al., 2023), an open-source dataset that integrates data from LLaVA-Instruct (Liu et al., 2024a) and COCO (Lin et al., 2014). Originally, ShareGPT4V contained 100,000 captions generated by GPT-4 Vision, later expanding to 1.2 million image-caption pairs using an improved captioning model while maintaining quality and scale.
To ensure diversity in synthetic data, we process data in batches. Each batch consists of 20,000 images and generates 5,000 synthetic conversations, with each conversation containing an average of 4.65 images. In total, we use 640,000 image-caption pairs to produce 160,000 synthetic conversations. We employ a vector sampling algorithm alongside long-form question prompts to ensure a diverse set of multi-step reasoning tasks. The modular design of our pipeline allows easy adaptation to new datasets and applications, making SMiR a scalable solution for multi-image reasoning.